Write the slope-intercept equation for the line that passes through (-2, 3) and is perpendicular to 2x – 9y = -1
Answer by mananth(16949)

2 x + -9 y = -1
Find the slope of this line

-9 y = -2 x -1
Divide by -9
y = 2/ 9 x + 1/9
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 2/9

The slope of a line perpendicular to the above line will be the negative reciprocal = -9/2
m1*m2=-1
m= -9/2 ,point (-2,3)
Find b by plugging the values of m & the point in
y=mx+b
3 = 9 + b
b= -6
m= -9/2
The required equation is y = -9/ 2 x-6
m.ananth@hotmail.ca